# Video2BEV: Transforming Drone Videos to BEVs for Video-based Geo-localization

### v1

@@ 6 Conclusion
In this work, we propose to leverage videos to mitigate the impact of environmental constraints in drone visual geo-localization. We propose a new Video2BEV paradigm that transforms drone-view videos into Bird’s Eye View (BEV) images by 3D gaussian splatting. This transformation effectively reduces the viewpoint gap between the drone view and the satellite view. Our Video2BEV paradigm also includes a diffusion-based module to generate negative samples, enhancing the scalability of the model. To validate the proposed method, we introduce the UniV dataset, a new video-based drone geo-localization dataset. The dataset includes flight paths of the drone at $30^{\circ}$ and $45^{\circ}$ elevation angles and corresponding videos recorded at up to 10 frames per second. Our Video2BEV paradigm outperforms other approaches in terms of Recall@1 and AP.
@@
Outline. This supplementary material includes two aspects:
@@
      • comparisons of the Video2BEV transformation at different elevation angles;
      • visualizations of drone-view videos, BEVs, and satellite-view images.
@@
    • more visualizations of synthetic negative samples.
@@ 7.1 Visualizations of the Video2BEV Transformation
Figure 9: Visualizations of drone-view videos, BEV videos, and satellite-view images. $i,j,k,m$ are frame indices, and $i<j<k<m$.
Figure 10: Elevation angles $\theta$ illustration in the UniV dataset. For each cases, top row shows $\theta$ = $45^{\circ}$ and bottom row displays $\theta$ = $30^{\circ}$. With a lower elevation angle, the new flight captures the target building with wider Field of View (FoV) but more occlusions, thereby posing more challenges for drone geo-localization.
Visualizations of Drone-view Videos, BEVs, and Satellite-view Images. We provide visualizations of images from different platforms. For each building, both drone-view and Bird’s Eye View (BEV) data are in video format and satellite-view data is in image format. Specifically, drones follow a spiral path around the target building, completing three circular flights. For BEVs, in the training set, we incorporate rotation angles and varying heights into BEV camera poses, generating a sequence of rotating and scaled-down BEVs (see Fig. 9). In the test set, we only increase the height of the BEV camera poses and render a sequence of scaled-down BEV images. The satellite view contains one image for each building. After the proposed Video2BEV transformation, the BEVs align with the same viewing direction as the satellite view and exhibit a similar color pattern to the drone view.
Figure 11: Visualizations of original images and synthetic hard negative samples.
@@ 7.2 Visualizations of the UniV Dataset
We provide additional visualizations of $45^{\circ}$ and $30^{\circ}$ elevation angles in the UniV dataset (see Fig. 10). Although both videos capture the same building, they differ significantly between the two elevation angles. Videos captured at a $45^{\circ}$ elevation angle provide overall views of the core areas of the target building, with these areas visible in most cases. In contrast, at the relatively lower elevation angle of $30^{\circ}$, drone-view videos offer a wider field of view but also introduce more occlusions. Consequently, core areas of the target building are occluded in some frames while remaining visible in others (see Fig. 10), effectively simulating outputs from real-world drone flights.
7.3 Visualizations of the Synthetic Negative Samples
@@ 8 Failure Case Analysis
We provide additional qualitative visualizations of retrieval results, with a particular focus on the failure cases (see Fig. 12). In these cases, the proposed method fails to recall the matched image in top-1. We observe that it is challenging because the recalled top-1 image has a very similar pattern to the query image, particularly in terms of the appearance and structure of the geographic target in the two images. In the first case, all recalled images share a similar structure and the predicted scores are relatively high. In the second case, all recalled images have a white rectangular roof. The roof of the ground truth image turns partly grey, which affects the retrieval prediction of our method. In the third case, recalled top-3 results have similar red roofs, making it challenging for the proposed method to accurately retrieve the ground truth building. In the last case, the recalled top-1 image has a similar architectural style to the query and the ground truth image is in shadow. Both factors contribute to an inaccurate retrieval result.
Figure 12: Typical failure cases for Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone. We observe that the failures are mainly due to two factors. First, some buildings were under construction, which is quite different from the current view. Second, some satellite-view photo color is not accurate, and some similar buildings are false-matched. Given queries (left) from different platforms, matched galleries are in green box, and mismatched galleries are in blue box. The scores on the top are similarity scores estimated by the proposed method.

### v3

@@ 6 Conclusion
In this work, we propose to leverage videos to mitigate the impact of environmental constraints in drone visual geo-localization. We propose a new Video2BEV paradigm that transforms drone-view videos into Bird’s Eye View (BEV) images by 3D gaussian splatting. This transformation effectively reduces the inter-platform viewpoint disparity between the drone view and the satellite view. Our Video2BEV paradigm also includes a diffusion-based module to generate negative samples, enhancing the intra-platform discriminative ability of the model. To support the video setting and validate the proposed framework, we introduce the UniV dataset, a new video-based drone geo-localization dataset. The dataset includes flight paths of the drone at $30^{\circ}$ and $45^{\circ}$ elevation angles and corresponding videos recorded at up to 10 frames per second. Extensive experiment validates that our Video2BEV paradigm outperforms other competitive approaches in both supervised setting on UniV and OOD testing on unseen SUES-200.
@@
Outline. This supplementary material includes 4 aspects:
@@
      • comparisons of the Video2BEV transformation at different elevation angles on the UniV;
      • visualizations of drone-view videos, BEVs, and satellite-view images on the UniV.
@@
    • more visualizations of synthetic negative samples on the UniV dataset.
@@
  3. 3.
    Implementation details.
  4. 4.
    Ablation study for loss weights.
@@ 7.1 Visualizations of the Video2BEV Transformation
Figure 8: The transformed BEV comparison of videos with different evaluation $\theta$ on the UniV. We highlight the challenging regions.
Figure 9: Visualizations of drone-view videos, BEV videos, and satellite-view images on the UniV dataset. $i,j,k,m$ are frame indices, and $i<j<k<m$.
Figure 10: Elevation angles $\theta$ illustration on the UniV dataset. For each case, top row shows $\theta$ = $45^{\circ}$ and bottom row displays $\theta$ = $30^{\circ}$. With a lower elevation angle, the new flight captures the target building with wider Field of View (FoV) but more occlusions, thereby posing more challenges for drone geo-localization.
@@
Visualizations of Drone-view Videos, BEVs, and Satellite-view Images. We provide visualizations of images from different platforms. For each building, both drone-view and Bird’s Eye View (BEV) data are in video format and satellite-view data is in image format. Specifically, drones follow a spiral path around the target building, completing three circular flights. For BEVs, in the training set, we incorporate rotation angles and varying heights into BEV camera poses, generating a sequence of rotating and scaled-down BEVs (see Fig. 9). The 2D rotation is designed for data augmentation. In the test set, we only increase the height of the BEV camera poses and render a sequence of scaled-down BEV images. The satellite view contains one image for each location. After the proposed Video2BEV transformation, the BEVs align with the same viewing direction as the satellite view and exhibit a similar color pattern to the drone view.
@@ 7.2 Visualizations of the UniV Dataset
We provide additional visualizations of $45^{\circ}$ and $30^{\circ}$ elevation angles in the UniV dataset (see Fig. 10). Although both videos capture the same building, they differ significantly between the two elevation angles. Videos captured at a $45^{\circ}$ elevation angle provide overall views of the core areas of the target building, with these areas visible in most cases. In contrast, at the relatively lower elevation angle of $30^{\circ}$, drone-view videos offer a wider field of view but also introduce more occlusions. Consequently, core areas of the target building are occluded in some frames while remaining visible in others (see Fig. 10), effectively simulating outputs from real-world drone flights. More visualizations can be found in the UniV-dataset.mp4.
7.3 Visualizations of Synthetic Negative Samples
Figure 11: Visualizations of original images and synthetic hard negative samples on the UniV dataset.
@@ 8 Failure Case Analysis
We provide additional qualitative visualizations of retrieval results, with a particular focus on the failure cases on the UniV (Fig. 12(a)) and SUES-200 (Fig. 12(b)) datasets. In these cases, the proposed method fails to recall the matched image in top-1. We observe that it is challenging because the recalled top-1 image has a very similar pattern to the query image, particularly in terms of the appearance and structure of the geographic target in the two images. In the first case, all recalled images share a similar structure and the predicted scores are relatively high. In the second case, all recalled images have a white rectangular roof. The roof of the ground truth image turns partly gray, which affects the retrieval prediction of our method. In the third case, recalled top-3 results have similar red roofs, making it challenging for the proposed method to accurately retrieve the ground truth building. In the fourth case, the recalled top-1 image has a similar architectural style to the query and the ground truth image is in shadow. Both factors contribute to an inaccurate retrieval result. In the last two cases, satellite-view photo color is not consistent with that of drone-view, resulting in false positive results.
Figure 12: Typical failure cases for Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone on the UniV (a) and SUES-200 (b) datasets. We observe that the failures are mainly due to two factors. First, some buildings were under construction, which is quite different from the current view. Second, some satellite-view photo color is not accurate, and some similar buildings are false-matched. Given queries (left) from different platforms, matched galleries are in green box, and mismatched galleries are in blue box. Scores on the top are similarity scores estimated from our method.
9 Implementation details
Data processing. (i) Video2BEV transformation. We estimate camera poses from drone-view videos using 8 NVIDIA GeForce RTX 4090 GPUs. The subsequent 3DGS training and BEV rendering are also conducted using the same computing resources. It takes less than 8 seconds to render 50 images with 512 $\times$ 512 with vanilla 3DGS. (ii) Hard Negative Sample Synthesis. It consists of caption generation, fine-tuning of the Stable Diffusion network, and hard sample synthesis, carried out on a single RTX 4090 GPU. All data will be released upon acceptance.
Model training. We adopt a two-stage training strategy. In the first stage, the encoder is based on the ViT architecture and supervised with instance loss and contrastive loss. In the second stage, the encoder is frozen and MLPs are optimized with the matching loss. All these experiments are conducted on 1 NVIDIA A800.
10 Ablation study for loss weights
Table 4: Ablation study for loss weights $\lambda$.
[…]
$\lambda$ R@1 AP R@1 AP
0.10 94.58 95.25 92.87 93.81
0.50 95.01 95.62 93.01 94.05
1.00 (Ours) 95.01 95.64 93.44 94.44
We employ a two-stage training strategy. In the first stage, the losses are the instance loss $\mathcal{L}_{I}$ and the contrastive loss $\mathcal{L}_{C}$, resulting in the total loss $\mathcal{L}=\mathcal{L}_{I}+\lambda\mathcal{L}_{C}$. The second stage utilizes the matching loss $\mathcal{L}_{M}$ only. By default, we set weights=1 following previous works. We add an ablation study on $\lambda$ in the first stage (Tab. 4). We observe that $\lambda=1$ achieves the best performance.